Joseph and Savannah are partners in a limited partnership. Joseph is the general partner with a 70% profits interest. Savannah is the limited partner with a 30% profits interest. Even though Savannah is a limited partner, she agreed to make an additional $20,000 capital contribution at any time the partnership required additional working capital. At the end of the year, the balance sheet showed $200,000 in recourse liabilities and an additional $60,000 in nonrecourse liabilities. Joseph’s beginning adjusted basis in his partnership interest was $80,000. Savannah’s beginning adjusted basis in her partnership interest was $50,000. Joseph’s and Savannah’s end-of-the year adjusted bases for their partnership interests would be:

A) Joseph $322,000, and Savannah’s $98,000.

B) Joseph $165,000, and Savannah’s $68,000.

C) Joseph $302,000, and Savannah’s $88,000.

D) Joseph $262,000, and Savannah’s $68,000.
Answer #1

A B 1 Ans: 2 C) Joseph S302,000, and Savannahs $88,000. 3 Particulars Joseph Savannahs 4 Opening $80,000 $50,000 5 $60,000А 1 Ans: 2 C) Joseph S 3 Particulars Joseph Savannahs 4 Opening 80000 50000 5 60000 =A5*70% =A5*30% 6 =200000-C6 20000 7 8 C